# Exercising the CCPA Opt-out Right on Android: Legally Mandated but Practically Challenging

### v4

@@ 3.3. Opting Out via GPC
The most important privacy preference signal for purposes of the CCPA is GPC. GPC is a draft standard at the W3C (Zimmeck et al., 2025). GPC signals are transmitted from users’ clients to servers via request headers. Servers can also query a user agent’s GPC status via the GPC DOM property. GPC can be set to 1 (header) or true (DOM property) to express a user’s opt-out as defined by applicable law in the jurisdiction of the user. Thus, GPC can be used in multiple jurisdiction and will have the meaning that a particular law or regulation attaches to it. GPC aims to enable users to exercise their opt-out rights beyond individual websites and apps for sets of sites or apps.
Recent studies evaluating the usability of GPC suggest that most users understand what it does and that they would turn it on if it were more broadly available (Zimmeck et al., 2023, 2024b). Its rate of adoption by website operators appears to increase slowly and steadily estimated at a few percentage points per year (Hausladen et al., 2025). Compliance with GPC is required in California (since January 2021) (Becerra, 2021), Colorado (since July 2024) (Colorado Department of Law, 2024), Connecticut (Office of the Attorney General, 2025) (since January 2025), and New Jersey (New Jersey Division of Consumer Affairs, 2025) (July 2025). The CCPA’s opt-out right and GPC apply not only to websites but also to mobile apps.
@@ 3.4. Opting Out via GPC on Mobile Apps
The Office of the California Attorney General brought various enforcement actions against mobile app operators for not honoring consumers’ opt-out right (State of California Department of Justice, 2022b, a, 2025). Consumers have a right to opt out on every “device,” which the CCPA defines as “any physical object that is capable of connecting to the internet, directly or indirectly, or to another device.” [^17] CCPA, §1798.140(o). Thus, the International Association of Privacy Professionals, a nonprofit and non-advocacy membership association of privacy professionals, recommends for mobile apps that “[a]t a minimum, mobile apps should […] include a functional ‘Do Not Sell or Share My Personal Information’ link in the app […]” (Desai and Zweifel-Keegan, 2023).
@@ 4.1. App Dataset
For our evaluation we downloaded a set of popular free apps from the US Google Play store in August 2023. Following previous work (Scoccia et al., 2020; Shuba and Markopoulou, 2020; Han et al., 2020), we identified the apps based on Google’s lists of popular apps in 49 app categories on the Google Play store. Each list contained 45 apps. The 49 app categories include 17 game categories (Google, 2024b). The sets of apps in the different categories are disjoint as each app is assigned exactly one category on the Google Play store (Google, 2024b). The 45 apps listed for each of the 49 categories are “heavily influenced by popularity” (Google, 2024a). [^19] Apps in the Communication and Productivity categories had the most installs while those in the Events category had the least. Appendix Figure 12 shows the app install ranges for the apps in all 49 app categories. We selected popular free apps for our study because those impact many users and are most likely to rely on an ad-based business model that implies selling personal information or sharing such with third parties. Using Raccoon (Onyxbits, 2022), we successfully downloaded 1,896/2,205 (86%) of the listed apps. The remaining apps were incompatible with our test device, removed from the Play Store before we could download them, or not successfully downloaded. We made two download attempts per app. For our GPC and AdID opt-out evaluation we successfully performed a dynamic analysis of 1,811/1,896 (96%) apps resulting in our app dataset.
@@ 4.2. Measuring Opt-out Effects
Figure 2. Exercising the CCPA opt-out right via GPC. 1. An HTTP request from the client Android device is sent with a GPC header. By sending such a header users are exercising their opt-out right per the CCPA. 2. HTTP responses contain scripts for setting privacy flags client-side. Identifying such privacy flags and their values is indicative for whether a user is opted out. App operators can implement the logic for processing privacy flags themselves or can use a CMP’s library. 3–6. Third-party ad networks can query the privacy flag via ad tags that the app operator integrates in its app. If the flag shows that the user has opted out, they are required to discard personal information such upon receipt.
[…]
Exercising opt-out rights via GPC on mobile apps.
@@
We aim to measure the effects of opting out—i.e., whether an app stopped selling or sharing personal information—in three ways: we measure whether an app (1) sends CCPA privacy flags set to opt-out to ad networks (such as Vungle’s ccpa status set to opted out (Vungle, 2025)), makes fewer connections to ad tracking domains, and (3) discloses fewer device identifiers. As to the first measure, it is the responsibility of third parties to provide functionality in their ad libraries for app operators to identify ad requests to which the CCPA applies. If an app operator enables such functionality, a privacy flag will be attached to all outgoing ad requests from the app to the third party. If a third party receives an ad request with a privacy flag, it knows to discard the targeting data upon receipt.
@@
Beyond changing privacy flag values, a successful opt-out can manifest itself through other observable changes in an app’s behavior. An app can comply by not making any ad calls. Thus, monitoring an app’s connections to third-party ad tracking domains serves as another indicator for opt-out adherence. Further, an app might connect to third-party ad domains but stop the sharing of identifying data. Thus, we also examine the impact of opt-out signals on data sharing practices, specifically, the disclosure of device identifiers. A reduction in the sharing of such identifiers can indicate that an opt-out request has been honored since a third party is now more limited in associating particular data with a particular user or device.
@@
Android’s ad ecosystem is integrated into the broader online ad ecosystem and relies on the delivery of ads via the web protocol. An opt-out request, and the corresponding setting of a privacy flag, can occur in several ways. Figure 2 shows how a user’s GPC opt-out request results in setting a privacy flag that is further propagated on the web, which is the underlying protocol for most ads on mobile apps. We evaluate to which extent such flags are set via in-app opt-out interfaces, GPC signals, and disabling the AdID. It is possible that some apps are not technically prepared to process GPC headers. Especially, since Android does not have a platform-level GPC setting, apps’ server-side code may well ignore GPC headers. This lack of explicit support for GPC, however, does not principally change the legal status of GPC as a mandatory opt-out setting on Android and, more generally, any web-based platform.
@@ 4.3. UI Opt-out Evaluation
For determining whether the CCPA opt-out right is applicable to an app we rely on app operators’ respective declarations in their privacy policies. For our app-level UI opt-out evaluation we first chose from our app dataset 100 random apps whose operators notify users in their privacy policies that they have a CCPA opt-out right. To ensure that a privacy policy is applicable to an analyzed app we relied on the link from each app’s US Google Play Store Data Safety page to the policy. If necessary, we further followed links in the policy to US state- or California-specific policies. The privacy policy analysis was performed by an author with legal expertise in California privacy law and attorney admitted to the California Bar Association. Similar to previous work (Kollnig et al., 2021), we categorized each app’s opt-out setting as follows:
@@
  • Opt-out Right Form: App has a CCPA right form, e.g., that requires users to enter their email address, to opt out of the selling or sharing of personal information or claim other CCPA rights.
  • Privacy Policy: App only has a privacy policy link but no opt-out setting inside the app. We did not follow any privacy policy-linked opt-out setting that is only accessible via the privacy policy. If there is just a link to a “privacy policy” inside the app, it would not be clear to a user that that is where they could opt out. Such opt-out setting would not be “clear and conspicuous” and insufficient per the CCPA if the app does not honor opt-out signals via GPC.
@@
We evaluated whether an app has an opt-out setting, for example, under a “privacy settings” or “settings” menu. If clicking on such a menu item took us to an external website, we followed the respective procedure there. For a subset of 38 apps—those with opt-out right setting or generic/SDK/cookie opt-out setting—we performed a dynamic analysis. We did not evaluate apps with opt-out right form because submitting such form would not have immediate impact visible in the apps’ network traffic. For 31/38 apps the network captures succeeded in our two test conditions. The two test conditions under which we evaluated each app are as follows:
@@
    Not Opted Out: Use of the main functionality of the app (e.g., play the game if the app is a game) with the app’s opt-out setting, if any, set to not opted out. This condition is the base condition under which we expect an app to sell and share users’ personal information per the CCPA.
@@
Figure 3. Overview of our dynamic analysis setup for capturing app network traffic. Appendix 10.1 contains a comprehensive description of our dynamic analysis for the UI and GPC and AdID opt-out evaluation.
[…]
Overview of our dynamic analysis setup for capturing app network traffic.
During our evaluation we performed for each app a manual dynamic analysis. For each condition, we captured each app’s network traffic for 30 seconds, similar to previous work (Kollnig et al., 2021), using a Google Pixel 6a with Android 13, our test device. We found in preliminary tests that this time period is sufficient to understand the app’s ad tracking behavior and that the order of the conditions does not impact the results. If necessary to use the app, we created an account and logged into the app before evaluating the two test conditions. For each app, when switching conditions, we cleared the persistent storage and shared preferences of the test device. Upon an app’s request, we granted all permissions except for location permission.
@@ 4.4. GPC and AdID Opt-out Evaluation
In addition to our app-level opt-out UI evaluation we performed a platform-level GPC and AdID opt-out evaluation. The “selling” or “sharing” of personal information by an app operator corresponds to a third party’s “buying” or “collecting.” Thus, to determine whether an app is selling or sharing personal information we identified the most prevalent third-party tracking domains across all apps in our app dataset based on the X-Ray Tracker List (Binns et al., 2018; Kollnig, 2023). [^20] We combined trackers with different subdomains in our app dataset into one domain to match the tracker domains in the X-Ray Tracker List. For example, we combined app.adjust.com, view.adjust.com, app.us.adjust.com, and s2s.adjust.com into adjust.com. We evaluated the 33 most prevalent third-party tracking domains across all apps in our app dataset to reach 99% coverage of at least one tracking domain in each app in our app dataset. We evaluated the privacy policies and developer documentation of the 33 third parties and looked for explicit references that—per the CCPA—they buy or collect personal information from app operators that integrate their libraries or sell or share such downstream.
@@
To illustrate, per InMobi’s privacy policy (InMobi, 2024), “InMobi may be deemed to be ‘selling’ personal information, […]”. In another example, Meta’s policy (Meta, 2024) indicates that Meta is buying or collecting personal information via its libraries by advising operators that they may “share with Meta Personal Information about California residents which would otherwise constitute a Sale or Sharing […].” We performed our evaluation conservatively. If a policy stated that a third party does not share or sell personal information per the CCPA, we accepted such statement as true.
@@
To evaluate the likelihood to which an app in our app dataset is part of a business, we selected a random sample of 100 apps from our app dataset for determining whether their operators self-declared to run a “business” in their privacy policies. We performed this analysis on the 83 sample apps in our app dataset that had a Google Play Data Safety page with a link to their privacy policy and were still available at the time of the sampling. If an app’s Data Safety page linked to a privacy policy, we assumed that the linked privacy policy applies to the app. The privacy policy analysis for our GPC and AdID opt-out evaluation was also performed by the author with privacy law expertise.
@@
The setup and procedure for our GPC and AdID opt-out evaluation is similar to our UI opt-out evaluation, though, automated. After downloading the apps, we ran an automatic dynamic analysis for each app individually proceeding in batches for a total of two weeks with the analysis of one app taking about 5 minutes. [^21] Appendix 10.1 contains further details of our dynamic analysis with Appendix Figure 10 showing the procedure in detail. We performed this analysis for the apps in our app dataset in October 2023. As shown in Figure 3, we instrumented the test device with a shell script, the controller, which ran on our laptop. The controller fetched an app from the app database on the laptop, installed it on the test device, and ran it under four test conditions.
@@
We want to find out if users can exercise their opt-out rights by sending GPC signals. We are also interested in the effect of disabling the AdID. Thus, we tested each app under four conditions encompassing all possible combinations of access to the AdID (yes/no) and sending GPC signals (yes/no). For each condition, we captured each app’s network traffic for 30 seconds. The four test conditions under which we evaluated each app in our app dataset are as follows:
@@
    AdID: The app has access to the AdID. This condition is the base condition under which we expect apps to sell and share users’ personal information per the CCPA.
@@
In the AdID + GPC and No AdID + GPC conditions we injected GPC headers into all HTTP requests using Mitmproxy. In the AdID and No AdID conditions, we injected dummy headers into all requests to minimize any differences that might occur due to header injection in the other two conditions. In all four conditions the controller granted all permissions requested by an analyzed app at install time. We tested whether granting or denying permissions made a difference for apps’ ad tracking behavior, which was not the case. [^22] Appendix Figure 13 shows the top 50 permissions that the 1,811 apps in our app dataset declared in their manifest file. When switching conditions, our controller cleared the persistent storage and shared preferences of the test device.

### v5

@@ 3.3. Opting Out via GPC
The most important privacy preference signal for purposes of the CCPA is GPC, which is a draft standard at the W3C (Zimmeck et al., 2025). GPC signals are transmitted from users’ clients to servers via request headers. Servers can also query a user agent’s GPC status via the GPC DOM property.
GPC can be set to 1 (header) or true (DOM property) to express a user’s opt-out as defined by applicable law in the jurisdiction of the user. Thus, GPC can be used in multiple jurisdiction and will have the meaning that a particular law or regulation attaches to it. GPC aims to enable users to exercise their opt-out rights beyond individual websites and apps for sets of sites or apps.
Recent studies evaluating the usability of GPC suggest that most users understand what it does and that they would turn it on if it were more broadly available (Zimmeck et al., 2023, 2024b). Its rate of adoption by website operators appears to increase slowly and steadily estimated at a few percentage points per year (Hausladen et al., 2025). Compliance with GPC is required in California (since January 2021) (Becerra, 2021), Colorado (since July 2024) (Colorado Department of Law, 2024), Connecticut (Office of the Attorney General, 2025) (since January 2025), and New Jersey (New Jersey Division of Consumer Affairs, 2025) (since July 2025). The CCPA’s opt-out right and GPC apply not only to websites but also to mobile apps.
@@ 3.4. Opting Out via GPC on Mobile Apps
The Office of the California Attorney General brought various enforcement actions against mobile app operators for not respecting consumers’ opt-out right (State of California Department of Justice, 2022b, a, 2025). Consumers have a right to opt out on every “device,” which the CCPA defines as “any physical object that is capable of connecting to the internet, directly or indirectly, or to another device.” [^17] CCPA, §1798.140(o). Thus, the International Association of Privacy Professionals, a nonprofit and non-advocacy membership association of privacy professionals, recommends for mobile apps that “[a]t a minimum, mobile apps should […] include a functional ‘Do Not Sell or Share My Personal Information’ link in the app […]” (Desai and Zweifel-Keegan, 2023).
@@ 4.1. App Dataset
For our evaluation we downloaded a set of popular free apps from the US Google Play store in August 2023. Following previous work (Scoccia et al., 2020; Shuba and Markopoulou, 2020; Han et al., 2020), we identified the apps based on Google’s lists of popular apps in 49 app categories on the Google Play store. Each list contained 45 apps. The 49 app categories include 17 game categories (Google, 2024b). The sets of apps in the different categories are disjoint as each app is assigned exactly one category on the Google Play store (Google, 2024b). The 45 apps listed for each of the 49 categories are “heavily influenced by popularity” (Google, 2024a). [^19] Apps in the Communication and Productivity categories had the most installs while those in the Events category had the least. Appendix Figure 12 shows the app install ranges for the apps in all 49 app categories. We selected popular free apps for our study because those are most likely to rely on an ad-based business model that implies selling personal information or sharing such with third parties (Section 3.1). Using Raccoon (Onyxbits, 2022), we successfully downloaded 1,896/2,205 (86%) of the listed apps. The remaining apps were incompatible with our test device, removed from the Play Store before we could download them, or not successfully downloaded. We made two download attempts per app. For our GPC opt-out evaluation we successfully performed a dynamic analysis of 1,811/1,896 (96%) apps resulting in our app dataset. For the remaining apps we did not collect any or only little network traffic, for example, due to apps’ crashing during the analysis.
@@ 4.2. Measuring Opt-out Effects
Figure 2. Exercising the opt-out right via GPC. 1. An HTTP request from the user’s Android device is sent with a GPC header. By sending this header the user is exercising the opt-out right. 2. The HTTP response contains a script for setting a privacy flag client-side. By identifying such a privacy flag and its value we determine whether the opt-out has been respected. App operators can implement the logic for processing privacy flags themselves or use the library of a CMP. 3–4. A third-party ad network queries the privacy flag via its ad tag that the app operator integrated in its app. 5. The flag may be attached to ad requests from the user to the ad network, which is the second point that we use to determine whether our opt-out has been respected. The ad network can also act directly upon receiving GPC signals, which are attached to all requests as headers. 6. The ad network responds to ad requests from opted out users with an ad not based on sold or shared personal information.:
@@
We aim to measure the effects of opting out—i.e., whether an app stopped selling or sharing personal information—in three ways: we measure whether an app (1) sends to ad networks CCPA privacy flags that are set to opt-out (such as Vungle’s ccpa status set to opted out (Vungle, 2025)), makes fewer connections to ad tracking domains, and (3) discloses fewer device identifiers. As to the first measure, it is the responsibility of ad networks to provide functionality in their ad libraries for app operators to identify ad requests to which the CCPA applies. If an app operator enables such functionality, a privacy flag should be set on the user’s device and attached to outgoing ad requests from the app to the third party, as shown in Figure 2. Thus, as privacy flags are generally communicated client-to-server, for example, ad networks retrieve the US Privacy String from an Android’s device local SharedPreferences (IAB Tech Lab, 2020), we can use client-side analysis to determine our opt-out status. If a third party receives an ad request with a privacy flag, it knows to discard personal information upon receipt and serve an ad not based on sold or shared personal information.
@@
Beyond changing privacy flag values, a successful opt-out can manifest itself through other observable changes in an app’s behavior. An app can comply by not making any ad calls. Thus, monitoring an app’s connections to third-party ad tracking domains serves as another indicator for opt-out adherence. Further, an app might connect to third-party ad domains but stop the sharing of identifying data. Thus, we also examine the impact of opt-out signals on data sharing practices, specifically, the disclosure of device identifiers. A reduction in the sharing of such identifiers can indicate that an opt-out request has been respected since a third party is now more limited in associating particular data with a particular user or device.
@@
Android’s ad ecosystem is integrated into the broader online ad ecosystem and relies on the delivery of ads via the web protocol. An opt-out request, and the corresponding setting of a privacy flag, can occur in several ways. Figure 2 shows how a user’s GPC opt-out request results in setting a privacy flag that is further propagated on the web, which is the underlying protocol for most ads on mobile apps. We evaluate to which extent such flags are set via in-app opt-out interfaces, GPC signals, and disabling the AdID. It is likely that many apps do not have functionality to process GPC headers. Especially, since Android does not have a platform-level GPC setting, apps operators may not have set up server-side code to capture GPC headers. This lack of explicit support for GPC, however, does not principally change the legal status of GPC as a mandatory opt-out setting on Android and, more generally, any web-based platform.
@@ 4.3. UI Opt-out Evaluation
For determining whether the CCPA opt-out right is applicable to an app we rely on app operators’ respective declarations in their privacy policies. For our app-level UI opt-out evaluation we first chose from our app dataset 100 random apps whose operators notify California residents in their privacy policies that they have a CCPA opt-out right. To ensure that a privacy policy is applicable to an analyzed app we relied on the link from each app’s US Google Play Store Data Safety page to the policy. If necessary, we further followed links in the policy to US state- or California-specific policies. The privacy policy analysis was performed by an author with legal expertise in California privacy law and attorney admitted to the California Bar Association. Similar to previous work (Kollnig et al., 2021), we categorized each app’s opt-out setting as follows:
@@
  • Opt-out Right Form: App has a CCPA right form, e.g., that requires California residents to enter their email address, to opt out of the selling or sharing of personal information or claim other CCPA rights.
  • Privacy Policy: App only has a privacy policy link but no opt-out setting inside the app. We did not follow any privacy policy-linked opt-out setting that is only accessible via the privacy policy. If there is just a link to a “privacy policy” inside the app, it would not be clear to a user that that is where they could opt out. Such opt-out setting would not be “clear and conspicuous” and insufficient per the CCPA if the app does not respect opt-out signals via GPC.
@@
We evaluated whether an app has an opt-out setting, for example, under a “privacy settings” or “settings” menu. If clicking on such a menu item took us to an external website, we followed the respective procedure there. For a subset of 38 apps—those with opt-out right setting or generic/SDK/cookie opt-out setting—we performed a dynamic analysis. We did not evaluate apps with opt-out right form because submitting such form would not have immediate impact visible in the apps’ network traffic. For 31/38 apps the network captures succeeded in our two test conditions. For the remaining apps we did not collect any or only little network traffic, for example, due to apps’ crashing during the analysis.
The two test conditions under which we evaluated each app are as follows:
@@
    Not Opted Out: Use of the main functionality of the app (e.g., if the app is a game, play it) with the app’s opt-out setting, if any, set to not opted out. This condition is the base condition under which we expect an app to sell and share California residents’ personal information.
@@
Figure 3. Overview of our dynamic analysis setup for capturing app network traffic. Appendix 10.1 contains a comprehensive description of our dynamic analysis for the UI and GPC opt-out evaluation.:
During our evaluation we performed for each app a manual dynamic analysis. For each condition, we captured each app’s network traffic for 30 seconds, similar to previous work (Kollnig et al., 2021), using a Google Pixel 6a with Android 13, our test device. We found in preliminary tests that this time period is sufficient to understand the app’s ad tracking behavior and that the order of the conditions does not impact the results. If necessary to use the app, we created an account and logged into the app before evaluating the two test conditions. For each app, when switching conditions, we cleared the persistent storage and SharedPreferences of the test device. Upon an app’s request, we granted all permissions except for location permission as an additional measure to prevent apps from detecting our real location.
@@ 4.4. GPC Opt-out Evaluation
In addition to our app-level opt-out UI evaluation we performed a platform-level GPC opt-out evaluation. The “selling” or “sharing” of personal information by an app operator corresponds to a third party’s “buying” or “collecting” of personal information. Thus, to determine whether an app is selling or sharing personal information we identified the most prevalent third-party tracking domains across all apps in our app dataset. Our starting point for this task was the X-Ray Tracker List (Binns et al., 2018; Kollnig, 2023). We found that 33 third-party tracking domains from the X-Ray Tracker List have a 99% coverage with at least one tracking domain in each app in our app dataset. [^20] We combined trackers with different subdomains in our app dataset into one domain to match the tracker domains in the X-Ray Tracker List. For example, we combined app.adjust.com, view.adjust.com, app.us.adjust.com, and s2s.adjust.com into adjust.com. For the 33 third parties associated with these domains we evaluated their privacy policies and developer documentation and looked for explicit references that—per the CCPA—they buy or collect personal information from app operators that integrate their libraries or sell or share such downstream.
@@
To illustrate, Meta’s policy (Meta, 2024) indicates that it is buying or collecting personal information per the CCPA via its libraries as it advises app operators integrating those that they may “share with Meta Personal Information about California residents which would otherwise constitute a Sale or Sharing […].” In another example, per InMobi’s privacy policy (InMobi, 2024), “InMobi may be deemed to be ‘selling’ personal information, […]” indicating InMobi’s further downstream sale of California residents’ personal information after buying or collecting such from the app operators who integrate its libraries. We performed our evaluation conservatively. If a policy stated that a third party does not share or sell personal information per the CCPA, we accepted such statement as true. Our methodology could be applied to cover cases where the operator claims not to sell or share personal information but does so in practice, for example, as it appears from the operator’s app network traffic, possibly in combination with a privacy flag not set to opt-out. In such case the operator needs to either adapt the claim or the practice.
@@
To evaluate the likelihood to which an app in our app dataset is part of a business, we selected a random sample of 100 apps from our app dataset for determining whether their operators self-declared to run a “business” in their privacy policies. We performed this analysis on the 83 sample apps in our app dataset that had a Google Play Data Safety page with a link to their privacy policy and were still available at the time of the sampling. If an app’s Data Safety page linked to a privacy policy, we assumed that the linked privacy policy applies to the app. The privacy policy analysis for our GPC opt-out evaluation was also performed by the author with privacy law expertise.
@@
The setup and procedure for our GPC opt-out evaluation is similar to our UI opt-out evaluation, though, automated. After downloading the apps, we ran an automatic dynamic analysis for each app individually proceeding in batches for a total of two weeks with the analysis of one app taking about 5 minutes. We performed this analysis for the apps in our app dataset in October 2023. As shown in Figure 3, we instrumented the test device with a shell script, the controller, which ran on our laptop. The controller fetched an app from the app database on the laptop, installed it on the test device, ran it in the foreground under each of the four test conditions, and uninstalled it. If an app update was triggered right after installing an app, it would not have been possible to run the app immediately afterwards, which we check for.  [^21] Appendix 10.1 contains further details of our dynamic analysis with Appendix Figure 10 showing the procedure in detail.
@@
We want to find out if California residents can exercise their opt-out rights by sending GPC signals. To which extent do Android apps respect GPC signals despite the lack of platform-level API support for such? We are also interested in the effect of disabling the AdID. Thus, we tested each app under four conditions encompassing all possible combinations of access to the AdID (yes/no) and sending GPC signals (yes/no). Similar to previous work (Kollnig et al., 2021), for each condition, we captured each app’s network traffic for 30 seconds. We also added a timeout of 30 seconds before running an app’s first condition so that the app is fully initialized before capturing of the network traffic begins (Nguyen et al., 2022). The four test conditions under which we evaluated each app in our app dataset are as follows:
@@
    AdID: The app has access to the AdID. This condition is the base condition under which we expect apps to sell and share California residents’ personal information.
@@
In the AdID + GPC and No AdID + GPC conditions we injected GPC headers into all requests, both HTTP and HTTPS requests, using Mitmproxy. In the AdID and No AdID conditions, we injected dummy headers into all requests to minimize any differences that might occur due to header injection in the other two conditions. In all four conditions the controller granted all permissions requested by an analyzed app at install time. [^22] Different from the UI Opt-out evaluation (Section 4.3), for the GPC opt-out evaluation we opted to grant apps location permission as we also wanted to explore location tracking. Appendix Figure 13 shows the top 50 permissions that the 1,811 apps in our app dataset declared in their manifest file. We tested whether granting or denying permissions made a difference for apps’ ad tracking behavior, which was not the case. When switching conditions, our controller cleared the persistent storage and SharedPreferences of the test device.